OpenStudy (hang254):

3c = 82 + c please state how you did it

OpenStudy (anonymous):

3c = 82 + c 2c = 82 c = 82 / 2 c = 41

OpenStudy (coolshubhs):

3c=82+c = 3c-c=82 2c=82 c=82/2 c=41

OpenStudy (hang254):

oh, forgot that c counted as 1

OpenStudy (cwrw238):

you want to solve for c i expect subtract c from both sides: 3c - c = 82 + c - c 2c = 82 now divide each side by 2 to get value of c

OpenStudy (hang254):

thanks to all

OpenStudy (cwrw238):

yw